OpenStudy (anonymous):

For the Function with rule f(x)=3x2^x+6, the average rate of change of f(x) between x=0 and x=2 is: help!

OpenStudy (anonymous):

The function is not defined properly. $$ f(x)=3x2^x+6$$ ??

OpenStudy (anonymous):

I think you just evaluate f(0), and f(2) and divide by 2

OpenStudy (anonymous):

its supposed to be 3 (times) 2^x + 6

OpenStudy (anonymous):

(f(2)-f(0)) / 2 should be the average rate of change

OpenStudy (anonymous):

how did u get that? i think you have to find the derivative first, i can do that, but im not sure what to do next

OpenStudy (anonymous):

the derivative gives you the instantaneous rate of change at any point, but the total change divided by the x-distance is the average rate of change

OpenStudy (anonymous):

Derivative is the instantaneous rate of change

OpenStudy (anonymous):

but (f(2)-f(0)) / 2= f'(c) by mean value theorem.

OpenStudy (anonymous):

with a line it's easy to see, e.g. f(x)=x, the average rate of change is 1 over any interval .. e.g. f(5)-f(2) / (5-2) = (5-2)/(5-2) = 1

OpenStudy (anonymous):

so you just want to do f(b) - f(a) / (b-a)

OpenStudy (anonymous):

the definition of the derivative is that quantity as b-a goes to 0

OpenStudy (anonymous):

Indeed.

OpenStudy (anonymous):

but if they're asking for the average over the entire interval you don't even have to differentiate, just compute f(b)-f(a) / (b-a)

OpenStudy (anonymous):

average rate of change = (f(x2)-f(x1))/x2-x1

OpenStudy (anonymous):

is that what your using?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

okay well the answer is 4.5, i dont see how 0-2 is going to give me that

OpenStudy (anonymous):

f(0) = 9 f(2) = 18 (18 - 9) / 2 = 9/2 = 4.5

OpenStudy (anonymous):

how did you get 9? isnt 2^0 = to 1?

OpenStudy (anonymous):

what is f(0)? plug 0 into your original equation everywhere that x appears, what do you get?

OpenStudy (anonymous):

ohh yepp, i see, i made a mistake. thanks!! i understand it

OpenStudy (anonymous):

ok